OpenStudy (anonymous):

What percentage of the area under a normal curve lies within the following boundaries: between the mean and a z of 1.27?

OpenStudy (anonymous):

Do you mean something like this?\[\frac{\int\limits_{0}^{1.27} e^{-x^2}dx}{\sqrt{\pi}}\]

OpenStudy (amistre64):

it depends on your table ..... if you are left tailed: z_0 - z_1.27 if your centered at the mean; then its just z_1.27 if its right tailed: z_1.27 - z_0

OpenStudy (amistre64):

if you have a calculator; normalCDF(0,1.27,0,1)